Asked to solve this question showing my work
The number of abortions worldwide decreased from 4
Show transcribed image text
Answer #1

Solved first four parts,post one more question to get the answer to remaining parts

A) Decrease in Number of abortion = Abortions in 1996 - Abortions in 2003

=> 46 million - 42 million

=> 4 million19

B) Number of abortion decrease per year = 4 million/(2003-1996) = 4/7 million = 571482.57

C) Decrease in Percentage = (46-42)/46 * 100 = 400/46 = 8.695%

d) Decrease in Percentage per year = 8.695/7 % = 1.2422%